ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 237, 2020

Can, M. B. and Y. Cherniavsky Stirling posets  185

Cherniavsky, Y. See M. B. Can

Corson, S. M. Automatic continuity of ℵ_{1}-free groups  267

D’Angeli, D., E. Rodaro and J. P. Wächter Automaton semigroups and groups: On the undecidability of problems related to freeness and finiteness

15

Dao, H. and I. Smirnov On the generalized Hilbert–Kunz function and multiplicity

155

Darji, U. B., M. Elekes, K. Kalina, V. Kiss and Z. Vidnyánszky The structure of random homeomorphisms

75

Delizia, C., U. Jezernik, P. Moravec and C. Nicotera Gaps in probabilities of satisfying some commutator-like identities

115

Dudko, A. and R. Grigorchuk On the question “Can one hear the shape of a group?” and a Hulanicki type theorem for graphs

53

Elekes, M. See U. B. Darji Farah, I., I. Hirshberg and A. Vignati The Calkin algebra is

ℵ_{1}-universal

287

Gekhtman, I., S. J. Taylor and G. Tiozzo Counting problems in graph products and relatively hyperbolic groups

311

Gitik, M. On κ-compact cardinals  457

Grigorchuk, R. See A. Dudko Grodzicki, W. Values of Iwahori-fixed vectors in the non-split Bessel

model on GSp(2n)

373

Hirshberg, I. See I. Farah Jezernik, U. See C. Delizia Kalina, K. See U. B. Darji Kiss, V. See U. B. Darji Klyachko, A. A. and M. A. Ryabtseva The dimension of solution

sets to systems of equations in algebraic groups

141

Maurischat, A. Reduced group schemes as iterative differential Galois groups

437

Moravec, P. See C. Delizia Nicotera, C. See C. Delizia

Ohsugi, H. and A. Tsuchiya Enriched chain polytopes  485

Praeger, C. E., J. Ramagge and G. A. Willis A graph-theoretic description of scale-multiplicative semigroups of automorphisms

221

Ramagge, J. See C. E. Praeger Reyes-Carocca, S. On Riemann surfaces of genus g with 4g − 4

automorphisms

415

Rodaro, E. See D. D’Angeli Ryabtseva, M. A. See A. A. Klyachko Smirnov, I. See H. Dao Taylor, S. J. See I. Gekhtman Tiozzo, G. See I. Gekhtman Tsuchiya, A. See H. Ohsugi Vidnyánszky, Z. See U. B. Darji Vignati, A. See I. Farah Wächter, J. P. See D. D’Angeli Willis, G. A. See C. E. Praeger Yang, Y. Arithmetical conditions of orbit sizes of linear groups of odd

order 1